### v1

# TokenSelect: Efficient Long-Context Inference and Length Extrapolation for LLMs via Dynamic Token-Level KV Cache Selection

Abstract.
With the development of large language models (LLMs), the ability to handle longer contexts has become a key capability for Web applications such as cross-document understanding and LLM-powered search systems. However, this progress faces two major challenges: performance degradation due to sequence lengths out-of-distribution, and excessively long inference times caused by the quadratic computational complexity of attention. These issues hinder the application of LLMs in long-context scenarios. In this paper, we propose Dynamic Token-Level KV Cache Selection (TokenSelect), a model-agnostic, training-free method for efficient and accurate long-context inference. TokenSelect builds upon the observation of non-contiguous attention sparsity, using Query-Key dot products to measure per-head KV Cache criticality at token-level. By per-head soft voting mechanism, TokenSelect selectively involves a small number of critical KV cache tokens in the attention calculation without sacrificing accuracy. To further accelerate TokenSelect, we designed the Selection Cache based on observations of consecutive Query similarity and implemented efficient dot product kernel, significantly reducing the overhead of token selection. A comprehensive evaluation of TokenSelect demonstrates up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end latency, while providing superior performance compared to state-of-the-art long-context inference methods.
Efficient LLMs Inference, Long-context LLMs, Sparse Attention
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the ACM on Web Conference; April 28–May 02, 2025; Sydney, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generation
1. Introduction
With the rapid development of large language models (LLMs), the number of parameters is no longer the sole factor significantly affecting model performance. The ability to effectively process longer context information has become one of the key metrics for evaluating LLMs’ capabilities. The latest Web applications such as cross-document understanding (Bai et al., 2024), LLM-powered search systems (Sharma et al., 2024), repository-level code completion (Zhang et al., 2023; Di et al., 2024), and complex reasoning (OpenAI, [n. d.]) have all placed higher demands on the long-context abilities of LLMs. There are two main difficulties in using pre-trained LLMs for long-context inference. On one hand, LLMs are limited by their context length during pre-training (e.g. Llama 3 only has 8192 tokens). Directly inferencing on longer sequences can lead to severe performance degradation due to reasons including sequence lengths out-of-distribution (Xiao et al., 2024; Han et al., 2024). On the other hand, even if LLMs possess sufficiently large context lengths, the quadratic computational complexity of attention with respect to sequence length makes the response time for long-context inference unbearable.
Figure 1. Distribution of tokens participating in attention computation under different sparsity patterns (indicated by blue dots). TokenSelect can more accurately select critical tokens for attention computation.
Previous works have made numerous attempts to address these difficulties. To extend the context length of LLMs, the current common practice is to perform post-training on long texts (Team et al., 2024; Yang et al., 2024; GLM et al., 2024). However, this approach comes with significant computational costs, particularly in two aspects: the synthesis of high-quality long-text data and the training process on extended sequences. To accelerate long-context inference, many studies focus on the sparsity of attention, attempting to reduce the scale of KV Cache involved in computation. The key to this type of method lies in designing sparse patterns for attention, which can be mainly divided into two categories: one uses predefined sparse patterns (Wang et al., 2019; Zaheer et al., 2020; Xiao et al., 2024; Han et al., 2024), while the other estimates the potential importance of KV Cache during the inference process (Zhang et al., 2024; Oren et al., 2024; Xiao et al., 2024; Tang et al., 2024; Jiang et al., 2024), attempting to select relevant KV Cache tokens into attention calculations. However, the design of these sparse patterns is often heuristically based on historical criticality or coarse-grained criticality estimation of tokens, making it difficult to ensure that the selected tokens are truly critical, thus resulting in sub-optimal performance, as shown in Fig. 1.
In this paper, we further observe the non-contiguous sparsity of attention, revealing the importance of designing more fine-grained dynamic sparse patterns. To this end, we propose TokenSelect, a model-agnostic and training-free approach that utilizes token-level selective sparse attention for efficient long-context inference and length extrapolation. Specifically, for each Query, TokenSelect dynamically calculates token-level per-head criticality for the past KV Cache and selects the $k$ most critical tokens through our head soft vote mechanism, involving them in the attention calculation. This reduces the scale of attention calculation to a constant length familiar to the model, while maintaining almost all of the long-context information, thereby simultaneously addressing the two main difficulties for long-context inference. To reduce the overhead of token selection, TokenSelect manages the KV Cache in token-level pages (Zheng et al., 2024) and design efficient kernel for token selection based on Paged KV Cache management through Triton (Tillet et al., 2019). Furthermore, based on our observation of high similarity between consecutive queries, we have designed the Selection Cache, which allows consecutive similar queries to share token selection results, thereby reducing the selection frequency while ensuring its effectiveness.
We evaluate the performance and efficiency of TokenSelect on three representative long-context benchmarks (Zhang et al., 2024; Hsieh et al., 2024; Bai et al., 2024) using three open-source LLMs (Yang et al., 2024; Dubey et al., 2024; AI et al., 2024). The experimental results demonstrate that our TokenSelect can achieve up to 23.84×23.84$\times$23.84 × speedup in attention computation compared to FlashInfer (flashinfer ai, [n. d.]), and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end inference latency compared to state-of-the-art long-context inference method (Xiao et al., 2024). Simultaneously, it provides superior performance on three long-text benchmarks. In summary, we make the following contributions:
  • An observation on the non-contiguous sparsity of attention that highlights the importance of token-level selection.
  • TokenSelect, a model-agnostic and training-free method that achieves accurate and efficient long-context inference and length extrapolation, which is compatible with mainstream LLM serving systems and ready for Web applications.
  • A comprehensive evaluation of TokenSelect, demonstrating up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end latency while exhibiting superior performance.
2. Related Works
Long-context LLMs
Due to computational complexity constraints, current LLMs based on Transformers often utilize limited context lengths during pre-training (Touvron et al., 2023; Dubey et al., 2024; Jiang et al., 2023; Yang et al., 2024; GLM et al., 2024; AI et al., 2024). To extend the long-context capabilities of LLMs, current methods can be broadly categorized into three approaches (Huang et al., 2024; Zhou et al., 2024; Zhao et al., 2024): 1) Modifying positional encodings: A widely adopted method is positional interpolation (Chen et al., 2023).  Chen et al. first proposed linear scaling of RoPE (Su et al., 2024) to map longer positional ranges within the original training window. Subsequent works (bloc97, 2023; emozilla, 2023) further improved this method using Neural Tangent Kernel (NTK) theory (Jacot et al., 2018), achieving longer context windows while maintaining model performance. Methods like YaRN (Peng et al., 2024) and Giraffe (Pal et al., 2023) optimize interpolation effects by adjusting frequency components or introducing temperature parameters. 2) Long-context post-training: This approach extends the model’s context length through additional training steps on longer documents after pre-training (Yang et al., 2024; Tian et al., 2024). It has been widely adopted by leading LLMs (Team et al., 2024; Yang et al., 2024; GLM et al., 2024) with the support of sequence parallelism techniques (Shoeybi et al., 2020; Jacobs et al., 2023; Liu et al., 2024). 3) Incorporating additional memory modules: Notable examples include Transformer-XL (Dai* et al., 2019), Compressive Transformer (Rae et al., 2020), RMT (Bulatov et al., 2022) and Infini-attention (Munkhdalai et al., 2024). Although these methods have expanded the context length of LLMs, long-context inference still faces the challenge of high computational costs.
Efficient Long-context Inference
In state-of-the-art LLMs serving systems (Kwon et al., 2023; Huggingface, 2024; NVIDIA, 2024; Zheng et al., 2024), technologies such as Flash Attention (Dao et al., 2022; Dao, 2024) and Paged Attention (Kwon et al., 2023) have greatly optimized LLMs inference efficiency by improving GPU I/O bottlenecks. However, in long-context inference scenarios, the quadratic computational complexity of attention with respect to sequence length poses new challenges for LLMs inference. Numerous studies focus on the sparsity of attention, selecting partial KV Cache for attention calculations to improve long-context inference efficiency. Sliding window (Wang et al., 2019; Zaheer et al., 2020) is one of the most widely used sparse patterns, reducing complexity to linear by executing attention computations within localized windows. Recent works like StreamingLLM (Xiao et al., 2024) and LM-infinite (Han et al., 2024) retain the initial tokens of the sequence in addition to sliding windows, effectively maintaining LLMs’ performance when processing long sequences. While these approaches are simple to implement, they cannot retain information from long contexts. Another approach focuses on dynamic KV Cache selection during inference. Methods like H2O (Zhang et al., 2024), TOVA (Oren et al., 2024), FastGen (Ge et al., 2024), Scissorhands (Liu et al., 2023), and SnapKV (Li et al., 2024) evaluate token criticality based on historical attention scores, selecting tokens within a limited budget. However, these methods permanently discard parts of the KV Cache, causing information loss from long contexts. To address this, InfLLM (Xiao et al., 2024) introduces Block Memory Units for KV Cache management, retrieving information from long contexts and offloading less-used blocks to CPU. Similarly, QUEST (Tang et al., 2024) proposes query-aware sparsity at page granularity, while MInference (Jiang et al., 2024) optimizes long-context inference using three sparse patterns. Apart from considering all attention heads, some other works (Ribar et al., 2024; Lee et al., 2024; Tang et al., 2024) attempt to focus on only a subset of attention heads. Beyond selection, some other research focuses on KV Cache quantization (Liu et al., 2024; Yang et al., 2024; He et al., 2024; Kang et al., 2024) and merging (Liu et al., 2024; Wan et al., 2024; Zhang et al., 2024; Wang et al., 2024). However, existing methods struggle to be applied in real-world Web applications, both in terms of accuracy and efficiency.
3. Preliminaries
In this section, we first introduce the inference process of LLMs, and then define the Selective Sparse Attention Problem.
3.1. LLMs Inference
Nowadays, mainstream LLMs are primarily based on the Decoder-only Transformer architecture, consisting sequentially of a word embedding layer, a series of transformer layers, and a token prediction head. Each transformer layer includes a multi-head attention (MHA) module and a feed-forward networks (FFN) module. The inference process of LLMs can be divided into two stages: the Prefill Stage and the Decode Stage.
The Prefill Stage is the preparatory phase of the inference process. In this stage, the user’s input is processed layer by layer through a single forward pass of LLMs, generating KV Cache for each layer. The generation of KV Cache is completed by the MHA module. Assuming $\mathbf{X}_{\text{prefill}}\in\mathbb{R}^{n_{\text{in}}\times d}$ is the input of a transformer layer, where $n_{\text{in}}$ is the number of tokens in user’s input sequence and $d$ is the hidden size. The computation of MHA in the Prefill Stage is as follows (simplified to single-head mode):
(1) $[\mathbf{Q}_{\text{prefill}},\mathbf{K}_{\text{prefill}},\mathbf{V}_{\text{prefill}}]=\mathbf{X}_{\text{prefill}}\cdot\left[\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\right],$
(2) $\mathbf{O}_{\text{prefill}}=\text{softmax}\left(\frac{\mathbf{Q}_{\text{prefill}}\cdot{\mathbf{K}_{\text{prefill}}}^{\top}}{\sqrt{d}}\right)\cdot\mathbf{V}_{\text{prefill}},$
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are linear projections, []delimited-[][$\cdot$][ ⋅ ] represents tensor concatenation operation, and Eq.(2) is also known as Scaled Dot-Product Attention (SDPA). After these computation, $\mathbf{K}_{\text{prefill}}$ and $\mathbf{V}_{\text{prefill}}$ are stored as the KV Cache for current layer $\mathbf{K}_{\text{cache}}$ and $\mathbf{V}_{\text{cache}}$, and $\mathbf{O}_{\text{prefill}}$ is used for subsequent calculations.
The Decode Stage is the phase where LLMs actually generate the response. In the Decode Stage, LLMs load the KV Cache and generate $n_{\text{out}}$ output tokens autoregressively through $n_{\text{out}}$ forward passes. Assuming $\mathbf{X}_{\text{decode}}\in\mathbb{R}^{1\times d}$ is the input of a transformer layer in a forward pass, the computation of MHA in the Decode Stage is as follows (The calculation of $\mathbf{Q}_{\text{prefill}}$ and $\mathbf{O}_{\text{prefill}}$ is consistent with that in the Prefill Stage):
(3) $\mathbf{K}_{\text{decode}}$ $=\left[\mathbf{K}_{\text{cache}},\ \mathbf{X}_{\text{decode}}\cdot\mathbf{W}_{k}\right],\;\mathbf{K}_{\text{cache}}\leftarrow\mathbf{K}_{\text{decode}},$
$\mathbf{V}_{\text{decode}}$ $=\left[\mathbf{V}_{\text{cache}},\ \mathbf{X}_{\text{decode}}\cdot\mathbf{W}_{v}\right],\;\mathbf{V}_{\text{cache}}\leftarrow\mathbf{V}_{\text{decode}},$
where $\mathbf{K}_{\text{decode}},\mathbf{V}_{\text{decode}}$ are composed of the KV Cache and the KV corresponding to the current input, which are then used to update the KV Cache of the current layer for use in the next forward pass.
LLMs inference, unlike training, is memory-bound, necessitating frequent GPU I/O operations between HBM and SRAM while underutilizing processing units. This bottleneck is particularly evident in SDPA computation. Optimizing for I/O is crucial for enhancing LLMs inference efficiency, especially in long-context scenarios.
3.2. Selective Sparse Attention
As discussed in the Sec. 1, the high attention sparsity in LLMs suggests sparse attention as a promising solution for long-context inference challenges. Sparse attention can keep the number of tokens participating in attention computations at a constant scale, rather than increasing with sequence length. Given that predefined sparse patterns are detrimental to performance, we aim to dynamically select crucial tokens for attention computation at each step during the inference process. Therefore, we formalize this problem according to the following definition.
(a) Attention is sparse in token-level.
(b) Block-level selection is sub-optimal.
(c) Attention logits is head-distinctive.
Figure 2. Motivations for token-level selection. (a) Visualization of attention scores sparsity. (b) Attention scores and critical token recalled by 1K token budget. (c) The $L_{1}$ norm of attention logits in each attention head. Visualizations are based on Qwen-2-7B-Instruct on the GovReport dataset of LongBench.
(a) Consecutive query is similar and the similarity distribution is consistent across datasets.
(b) Selection overlaps when query is similar.
Figure 3. Observations on similarity of consecutive queries. (a) Cosine similarity distribution between consecutive queries. (b) The token selection overlap rate ($\frac{|\mathcal{I}_{i}\cap\mathcal{I}_{i+1}|}{|\mathcal{I}_{i+1}|}$) with respect to consecutive query similarity.
Definition 1 (Selective Sparse Attention Problem, informal). For current input of length $C$ ($C=1$ in the Decode Stage) and KV Cache of length $N$, assuming there are $H$ attention heads with a head size of $d_{h}$, let $\mathbf{O}$ be the output of the SDPA:
(4) $\mathbf{O}=\left[\text{softmax}\left(\frac{\mathbf{Q}^{h}\cdot{\left[\mathbf{K}^{h}_{\text{cache}},\ \mathbf{K}^{h}_{\text{current}}\right]}^{\top}}{\sqrt{d}}\right)\cdot[\mathbf{V}^{h}_{\text{cache}},\ \mathbf{V}^{h}_{\text{current}}]\right]_{h\in\{1,\cdots,H\}},$
where $\mathbf{Q}^{h},\mathbf{K}^{h}_{\text{current}},\mathbf{V}^{h}_{\text{current}}\in\mathbb{R}^{C\times d_{h}}$ is the Query, Key, Value matrices of current input for head $h$ and $\mathbf{K}^{h}_{\text{cache}},\mathbf{V}^{h}_{\text{cache}}\in\mathbb{R}^{N\times d_{h}}$ represent the KV Cache. Let $\hat{\mathbf{O}}$ be the output of the Selective Sparse Attention:
(5) $\hat{\mathbf{O}}=\left[\text{softmax}\left(\frac{\mathbf{Q}^{h}\cdot{\left[\mathbf{K}^{h}_{\text{select}},\ \mathbf{K}^{h}_{\text{current}}\right]}^{\top}}{\sqrt{d}}\right)\cdot[\mathbf{V}^{h}_{\text{select}},\ \mathbf{V}^{h}_{\text{current}}]\right]_{h\in\{1,\cdots,H\}},$
where $\mathbf{K}^{h}_{\text{select}},\mathbf{V}^{h}_{\text{select}}\in\mathbb{R}^{k\times d_{h}}$ are $k$ selected KV Cache ($k\ll N$). The selection of $\mathbf{K}_{\text{select}},\mathbf{V}_{\text{select}}$ is performed by selection function $\mathcal{S}$:
(6) $\mathcal{S}\left(\mathbf{Q},\ \mathbf{K}_{\text{cache}}\right)=\mathcal{I},\;\text{where}\;\mathcal{I}\in\mathcal{P}(\{1,\cdots,N\}),$
$\mathbf{K}_{\text{select}}=[(\mathbf{K}_{\text{cache}})_{i}]_{i\in\mathcal{I}},\;\mathbf{V}_{\text{select}}=[(\mathbf{V}_{\text{cache}})_{i}]_{i\in\mathcal{I}},$
where $\mathcal{I}$ is the set of selected indices. The objective is to find an appropriate selection function $\mathcal{S}$ that minimizes the difference between the outputs of the SDPA and the selective sparse attention:
(7) $\min_{S}\left\|\mathbf{O}-\hat{\mathbf{O}}\right\|_{2}^{2}.$
Existing works on long-context inference (Wang et al., 2019; Zaheer et al., 2020; Xiao et al., 2024; Han et al., 2024; Zhang et al., 2024; Oren et al., 2024; Li et al., 2024; Xiao et al., 2024; Tang et al., 2024; Jiang et al., 2024) can be categorized under the Selective Sparse Attention Problem, with variations primarily in the design of the selection function $\mathcal{S}$.  (Wang et al., 2019; Zaheer et al., 2020; Xiao et al., 2024) have developed input-independent selection functions $\mathcal{S}()$, while  (Zhang et al., 2024; Oren et al., 2024; Li et al., 2024) propose query-independent functions $\mathcal{S}(\mathbf{K}_{\text{cache}})$ for improved performance. Current state-of-the-art methods (Xiao et al., 2024; Tang et al., 2024; Jiang et al., 2024) utilize query-aware selection functions $\mathcal{S}(\mathbf{Q},\mathbf{K}_{\text{cache}})$. However, these approaches typically operate at a block-level, which limits their effectiveness and overall performance.
4. Motivations and Observations
Attention is Sparse, Non-contiguous and Head-Distinctive
Previous works (Wang et al., 2019; Zaheer et al., 2020; Xiao et al., 2024; Han et al., 2024; Zhang et al., 2024; Oren et al., 2024; Li et al., 2024; Xiao et al., 2024; Tang et al., 2024; Jiang et al., 2024) have demonstrated the sparsity of attention scores in LLMs, particularly when processing long texts. Recent approaches (Xiao et al., 2024; Jiang et al., 2024; Tang et al., 2024) partition the KV Cache into non-overlapping blocks, estimating block criticality for sparse attention calculations. These methods assume that tokens with higher attention scores tend to be contiguous. However, our further observations reveal that this assumption does not always hold true in practice. As illustrated in Fig. 2a, attention scores are sparsely distributed at the token-level, with critical tokens not necessarily contiguous. This non-contiguity leads to significant omissions in block-level token selection. Fig. 2b demonstrates that finer selection granularity improves recall of critical tokens, motivating us to perform token-level selection. For token-level selection, an intuitive approach would be to directly select the top-$k$ tokens with the highest attention logits. However, observation in Fig. 2c reveals considerable disparity in the $L_{1}$ norm of attention logits across attention heads. As a result, the selection result tends to be dominated by a few heads with disproportionately large attention logits, driving us to design a more robust selection function that maintains the independence of heads.
Consecutive Queries are similar
As sparsity of attention is dynamic (Zhang et al., 2024; Xiao et al., 2024; Tang et al., 2024; Jiang et al., 2024), token selection should be performed for every Query, which inevitably increases the computational overhead of selective sparse attention. Fortunately, we observe that consecutive Queries exhibit high similarity, as shown in Fig. 3a. Intuitively, when two consecutive Queries are highly similar, their dot products with the Keys will also be similar, leading to substantial overlap in the token selection results. Due to space constraints, we provide an informal lemma about this below. The formal version and corresponding proof can be found in the Appendix A.
Figure 4. The illustration of TokenSelect, which involves calculating per-head criticality using the Paged Dot Product Kernel, performing head soft vote to get selected indices, and executing selective sparse attention via the Paged Attention Kernel.
Lemma 1 (Informal). Consider Queries $\mathbf{Q}_{1},\mathbf{Q}_{2}\in\mathbb{R}^{1\times d}$ that are consecutive and a Key set $\{\mathbf{K}_{i}\}_{i=1}^{N}$. Let $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ be the sets of indices of the top-$k$ Keys selected by dot product for $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$ respectively. If $cos(\mathbf{Q}_{1},\mathbf{Q}_{2})>\epsilon$, where $\epsilon$ is a threshold, then $\mathcal{I}_{1}=\mathcal{I}_{2}$.
Fig. 3b illustrates this lemma experimentally. It can be seen that the overlap rate of token selection tends to increase with query similarity. This key insight motivates us to reuse selection results for similar queries, improving computational efficiency. Moreover, the similarity distribution of consecutive Queries remains consistent across different tasks, as demonstrated in Fig. 3a, allowing us to apply a global similarity threshold across all scenarios.
5. Designs of TokenSelect
In this section, we will introduce the design details of TokenSelect, primarily encompassing the Selection Function, the Selection Cache, and efficient implementation of TokenSelect. The overall workflow of TokenSelect is illustrated in Fig. 4.
5.1. Selection Function
The simplest selection function is to determine the criticality of the tokens through the dot product of $\mathbf{Q}$ and $\mathbf{K}_{\text{cache}}$, then select the top-$k$ critical ones as $\mathbf{K}_{\text{select}},\mathbf{V}_{\text{select}}$. The selected indices $\mathcal{I}$ are calculated as follow:
(8) $\mathcal{I}_{\text{topk}}=\operatorname{TopK}\left(\mathbf{Q}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right).$
However, as discussed in Sec. 4, this approach is prone to inaccuracies due to disparities in norm of attention logits between heads. To maintain independence between heads, a better approach is to have each head select the top-$k$ most critical tokens, and then determine the final selection through voting among the heads:
(9) $\mathcal{I}_{\text{head-vote}}=\operatorname{TopK}\left(\sum_{h=1}^{H}\mathbb{I}\left(i\in\operatorname{TopK}\left(\mathbf{Q}^{h}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right)\right)\right),$
where $\mathbb{I}$ is the indicator function. Unfortunately, despite better performance, this method relies on scatter_add and multiple topk operations, resulting in low efficiency on GPUs. Additionally, the 0/1 voting ignores the relative importance of tokens for each head. Therefore, we propose a head soft vote approach that offers better performance and efficiency. Specifically, we first calculate the per-head criticality, then normalize through softmax, and sum the results for all heads. The formalization is as follows:
(10) $\mathcal{I}_{\text{head-soft-vote}}=\operatorname{TopK}\left(\sum_{h=1}^{H}\operatorname{softmax}\left(\mathbf{Q}^{h}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right)\right).$
5.2. Optimizing Selection Frequency
Although the aforementioned selection function can reduce the complexity of attention from $O(N^{2})$ to $O(k^{2})$ ($k\ll N$), while maintaining performance, the execution time of the selection function itself still affects the latency of LLMs inference. To further accelerate long-context inference, based on our observations of the similarity of consecutive queries, we design optimization strategies for both the Prefill Stage and the Decode Stage to reduce the selection frequency while ensuring its effectiveness.
In the Prefill Stage, $\mathbf{Q}_{\text{prefill}}\in\mathbb{R}^{n_{\text{in}}\times d}$ is inputed. In long-context scenarios, the number of tokens in the user’s input sequence $n_{\text{in}}$ may reach up to 1M, making it impractical to perform selection for each Query token. Considering the similarity of consecutive Queries, we use chunk-wise token selection, inputting $\frac{1}{c}\sum_{i=1}^{c}(\mathbf{Q}_{C})_{i}$ into the selection function, where $\mathbf{Q}_{C}\in\mathbb{R}^{c\times d}$ is the Query chunk and $c$ is the chunk size. This method helps maintain the compute-intensive nature of the Prefill Stage, preventing it from becoming memory bound.
Algorithm 1 Selection Cache Algorithm for the Decode Stage
0:  $\mathbf{Q}$: current Query,  $k$: number of selected tokens,      $\mathbf{C}_{Q}$: Query cache, $\mathbf{C}_{\mathcal{I}}$: selection cache,      $\mathcal{S}$: selection function (Eq.(10)), $\theta$: similarity threshold,      $\mathbf{f}$: first query flags (default True)
0:  $\mathcal{I}$: selected indices
1:  if $\mathbf{f}$ or $\text{cos}(\mathbf{Q},\mathbf{C}_{Q})<\theta$ then
2:     $\mathcal{I}\leftarrow\mathcal{S}(\mathbf{Q},k)$
3:     $\mathbf{C}_{\mathcal{I}}\leftarrow\mathcal{I}$
4:     $\mathbf{C}_{Q}\leftarrow\mathbf{Q}$
5:     $\mathbf{f}\leftarrow\text{False}$
6:  else
7:     $\mathcal{I}\leftarrow C_{\mathcal{I}}$
8:  end if
9:  return  $\mathcal{I}$
In the Decode Stage, due to the auto-regressive characteristic of LLMs, we need to frequently perform selection for $\mathbf{Q}_{\text{decode}}\in\mathbb{R}^{1\times d}$, and this process cannot be executed chunk-wise like in the Prefill Stage. To reduce the frequency of token selection in the Decode Stage, as shown in Algorithm 1, we propose the Selection Cache. Consecutive similar Queries will hit the cache, thereby directly loading the cached selection results for the previous Query. The Selection Cache allows us to reduce decode latency while maintaining almost the same performance.
Table 1. Comparison of different methods with different origin models on InfiniteBench.
Methods En.Sum En.QA En.MC En.Dia Code.D Math.F R.PK R.Num R.KV Avg.
Qwen2-7B 23.80 14.92 54.59 8.50 28.17 19.71 28.81 28.64 19.00 25.13
NTK 18.73 15.34 41.28 7.50 24.87 27.71 99.15 97.46 59.80 43.54
SelfExtend 3.76 4.44 20.09 5.00 8.12 2.29 0.00 0.00 0.00 4.86
StreamingLLM 19.60 13.61 48.03 3.50 27.92 19.43 5.08 5.08 2.40 16.07
InfLLM 19.65 15.71 46.29 7.50 27.41 24.00 70.34 72.20 5.40 32.06
TokenSelect 22.62 18.86 48.47 7.50 30.20 32.57 100.00 100.00 86.60 49.65
Llama-3-8B 24.70 15.50 44.10 7.50 27.92 21.70 8.50 7.80 6.20 18.21
NTK 6.40 0.40 0.00 0.00 0.50 2.60 0.00 0.00 0.00 1.10
SelfExtend 14.70 8.60 19.70 0.00 0.00 22.60 100.00 100.00 0.20 29.53
StreamingLLM 20.40 14.30 40.60 5.00 28.43 21.40 8.50 8.30 0.40 16.37
InfLLM 24.30 19.50 43.70 10.50 27.41 23.70 100.00 99.00 5.00 39.23
TokenSelect 26.99 19.39 45.85 14.50 27.41 28.29 100.00 97.29 40.00 44.41
Yi-1.5-6B 18.78 10.48 39.74 5.00 29.95 16.00 5.08 5.08 0.00 14.45
NTK 4.66 0.58 0.87 0.00 0.00 1.43 0.00 0.00 0.00 0.83
SelfExtend 5.62 1.07 1.31 0.00 0.00 1.14 0.00 0.00 0.00 1.01
StreamingLLM 15.35 9.26 35.81 5.00 27.41 14.29 5.08 4.92 0.00 13.01
InfLLM 16.98 8.93 34.06 3.00 27.41 16.86 100.00 96.61 0.00 33.76
TokenSelect 21.13 12.32 40.61 5.50 30.71 20.86 100.00 99.83 0.00 36.77
5.3. Efficient Implementation
To ensure that our proposed TokenSelect can be used for real-world Web applications, efficient implementation is crucial. We first analyze the computation time breakdown of representative block-level selective sparse attention method, InfLLM (Xiao et al., 2024). From (1)(2)(3) in Fig. 5, we can observe that although selective sparse attention can significantly reduce the complexity of attention calculations, the actual computation time is still highly dependent on the implementation. The incompatibility with efficient attention implementations such as Flash Attention has resulted in methods requiring historical attention scores (Zhang et al., 2024; Oren et al., 2024; Li et al., 2024; Xiao et al., 2024) being difficult to be applied in real-world Web applications. Through the analysis of InfLLM’s Flash Attention-compatible version, we make several discoveries. The initial motivation for estimating token criticality at the block-level is to reduce the overhead of selection function (mainly considering dot product calculation). However, we find that dot product is not the primary performance bottleneck. Instead, a significant portion of the overhead comes from indexing the KV Cache using selected indices and making them contiguous in GPU memory, which frequently occurs during the updating of KV blocks and the concatenation of selected KV Cache. The extensive I/O required for this operation further exacerbates the memory-bound in LLMs inference. Based on this, we propose that Paged Attention is a more suitable implementation for selective sparse attention. Using Paged KV Cache management (with page size=1 for TokenSelect), we can reduce the I/O volume for selection results from the scale of all selected KV Caches $O(2kd)$ to the scale of their indices $O(k)$. However, by observing (4) in Fig. 5, we find that we encounter another bottleneck under Paged KV Cache management. Since logically contiguous KV Cache is not entirely contiguous in GPU memory, it also needs to be made contiguous before performing computational operations. To address this issue, we draw inspiration from the concept of Paged Attention and implement a Paged Dot Product Kernel using Triton (Tillet et al., 2019), which significantly improves the overall efficiency of TokenSelect.
Figure 5. Computation time breakdown for single chunk prefill step under different attention implementations (chunk size: 512, KV Cache length: 128K, attended tokens: 4K).
6. Experiments
In this section, we first introduce the experimental setup of this paper, and then reveal the performance and efficiency of our TokenSelect in long-context inference through experiments.
6.1. Experimental Settings
Datasets.
To evaluate TokenSelect’s performance on long-context inference, we use the following datasets: (1) InfiniteBench (Zhang et al., 2024): The mainstream long-context benchmark consisting of multi-tasks. The average length of it exceeds 200K tokens. (2) RULER (Hsieh et al., 2024): A challenging long-context benchmark containing 13 different tasks, with subsets of varying lengths up to 128K tokens. (3) LongBench (Bai et al., 2024): Another mainstream long-context benchmark comprising 6 types of tasks. The 95% percentile for its lengths is 31K tokens. For each dataset, we use its recommended metrics, which are presented in the Appendix B.
Baselines.
To demonstrate the state-of-the-art (SOTA) performance of TokenSelect, we include the following methods for comparison: (1) Original models: We select three mainstream open-source LLMs - Qwen2-7B-Instruct (Yang et al., 2024), Llama-3-8B-Instruct (Dubey et al., 2024), and Yi-1.5-6B-Chat (AI et al., 2024) - utilizing their original context lengths without any modifications. (2) NTK-Aware Scaled RoPE: A nonlinear RoPE interpolation method. (3) SelfExtend: A RoPE interpolation method that reuses the position ids across neighboring tokens. (4) StreamingLLM: The SOTA method for long-context inference with predefined sparse patterns. (5) InfLLM: The SOTA method for long-context inference and length extrapolation using a block-level selective sparse attention method. (6) MInference: The SOTA method for long-context prefilling acceleration, utilizing three sparse patterns including block-level sparse attention. It’s worth noting that since MInference doesn’t support length extrapolation, we use an alternative evaluation method, applying it to Llama-3-8B-Instruct-262k (Llama3 after long-text post-training). Additionally, we do not include another SOTA method, QUEST (Tang et al., 2024), as it does not support Grouped Query Attention (GQA).
Table 2. Comparison of different methods with different origin models on RULER.
Methods 4K 8K 16K 32K 64K 128K Avg.
Qwen2-7B 90.74 84.03 80.87 79.44 74.37 64.13 78.93
StreamingLLM 94.41 54.59 33.54 22.40 15.38 10.88 38.53
InfLLM (2K+512) 52.85 36.09 29.36 23.52 18.81 18.29 29.82
InfLLM (4K+4K) 55.22 52.10 40.53 29.77 21.56 18.64 36.30
Ours (2K+512) 94.11 81.81 68.68 60.62 51.81 42.75 66.63
Ours (4K+4K) 94.42 90.22 82.06 70.40 59.66 54.28 75.17
Llama-3-8B 93.79 90.23 0.09 0.00 0.00 0.00 30.69
StreamingLLM 93.68 54.48 33.77 20.35 14.88 11.47 38.11
InfLLM (2K+512) 79.79 52.43 40.12 33.60 25.68 23.39 42.50
InfLLM (4K+4K) 93.79 86.11 64.33 45.39 33.13 27.81 58.43
Ours (2K+512) 93.73 82.92 71.92 65.38 59.35 33.39 67.78
Ours (4K+4K) 93.88 90.29 70.13 57.72 48.36 39.38 66.63
Yi-1.5-6B 73.12 9.09 0.37 0.01 0.00 0.01 13.77
StreamingLLM 72.10 33.03 21.69 15.39 12.58 12.61 27.90
InfLLM (2K+512) 59.66 36.77 27.41 24.49 21.49 21.17 31.83
InfLLM (4K+4K) 74.81 52.57 27.65 22.83 20.19 19.48 36.26
Ours (2K+512) 75.93 59.55 49.69 42.36 34.68 31.36 48.93
Implementation details.
[…]
6.2. Performance Comparisons
InfiniteBench.
As shown in Table 1, our TokenSelect achieves significantly superior overall performance on InfiniteBench compared to all baseline methods, even though TokenSelect uses the smallest token budget (¡3K). The fact that it significantly outperforms the original models demonstrates TokenSelect’s strong length extrapolation capability. We analyze that this is due to our adoption of a fine-grained KV Cache selection strategy, while considering the equal contribution of each head to selection, which ensures that we can select most critical tokens. Observing the performance of other methods, we find that RoPE interpolation methods (NTK, SelfExtend) generally perform poorly unless used on specially trained models such as Qwen2-7B-Instruct. The better performance of Qwen2-7B-Instruct on the original model can also be attributed to this. The sparse attention method StreamingLLM, based on fixed sparse patterns, can guarantee some of the model’s capabilities, but due to discarding a large amount of long-context information, it performs poorly on retrieval-related tasks (R.PK, R.Num, R.KV). The block-level selection method InfLLM can retain more long-context information compared to StreamingLLM. However, due to its sub-optimal block-level selection, it results in lower performance on most tasks compared to TokenSelect, even though we set a larger token budget for InfLLM. It is worth noting that Yi-1.5-6B does not perform normally on the R.KV task, as it is unable to correctly recite strings like the Universally Unique Identifier.
(a) LongBench-GovReport.
(b) InfiniteBench-En.QA.
(c) InfiniteBench-R.PK.
(d) InfiniteBench-R.KV.
Figure 6. Performance and Cache Rate with different similarity threshold $\theta$ of the Selection Cache on Qwen2-7B-Instruct.
RULER
To further demonstrate the long-context capability of TokenSelect, we conduct evaluation on the more challenging long-context benchmark RULER. Considering the increased difficulty of RULER and its substantial computational requirements, we include only comparable baseline methods. As shown in Table 2, our TokenSelect maintains significantly superior overall performance compared to other long-context inference methods. For all models, TokenSelect achieves length extrapolation while preserving the model’s original capabilities, benefiting from our efficient utilization of the model’s limited context length. Notably, due to the constraints of model’s context length, TokenSelect experiences performance degradation with larger token budgets (4K+4K) on Llama and Yi. However, its performance with smaller token budgets still significantly surpasses other baseline methods.
LongBench.
Due to space constraints, the results of LongBench are presented in the Appendix C. Although its relatively shorter text length makes it less suitable for evaluating state-of-the-art long-context inference methods, our TokenSelect still demonstrates superior overall performance compared to most baseline methods.
Table 3. Comparison of different methods on post-trained models on InfiniteBench and LongBench.
Methods InfiniteBench LongBench
En.Sum En.QA Code.D Math.F R.KV Avg.
LLaMA-3-8B-262K 20.2 12.4 22.1 26.6 14.4 33.9
    + MInference 20.5 12.9 22.3 33.1 12.8 38.4
Ours (w/ LLaMA-8K) 26.9 19.3 27.4 28.2 40.0 44.0
Comparing to methods based-on post-trained model.
In Table 3, we present the performance of the post-trained model and long-context inference method (Jiang et al., 2024) based on it. It shows that even compared to length extrapolation methods requiring additional training, the training-free TokenSelect still exhibits superior performance on most tasks. Although Minference can improve the performance of the original model, it fails to reverse the negative impact of long-text post-training on shorter text tasks (LongBench).
Table 4. Ablation study of the Selection Function $\mathcal{S}$ on InfiniteBench using Qwen2-7B-Instruct.
$\mathcal{S}$ En.QA En.MC Code.D Math.F R.Num R.KV
$\mathcal{I}_{\text{topk}}$ 15.15 45.85 28.43 31.14 98.47 16.60
$\mathcal{I}_{\text{head-vote}}$ 17.01 45.85 28.68 30.86 100.00 22.40
$\mathcal{I}_{\text{head-soft-vote}}$ 18.86 48.47 30.20 32.57 100.00 86.60
6.3. Ablation Studies
In ablation studies, we primarily analyze the impact of different Selection Functions $\mathcal{S}$ on performance. To compare the performance of different Selection Functions $\mathcal{S}$ under low token budgets (i.e., token efficiency), we maintain the 2K+512 configuration. From Table 4, we can observe that our proposed head soft vote mechanism performs significantly better across all tasks. This indicates that using the head soft vote mechanism to balance each head’s contribution to token selection results can help us avoid the domination of selection by few heads with large attention logits.
Table 5. Performance vs. Number of selected tokens $k$ on InfiniteBench using Qwen2-7B-Instruct.
$k$ En.Sum En.QA En.Mc Math.F R.Num R.KV
128 21.23 10.46 41.48 18.00 100.00 13.40
256 22.01 11.66 41.92 19.71 100.00 20.00
512 21.60 13.31 40.17 21.71 100.00 45.60
1K 21.35 15.13 44.10 24.57 100.00 73.00
2K 22.62 18.86 48.47 32.57 100.00 86.60
4K 24.09 21.11 51.53 21.71 100.00 88.00
8K 25.32 22.93 58.52 23.71 100.00 85.40
16K 26.54 23.04 62.88 28.16 100.00 72.00
(a) KV Cache length: 128K.
(b) KV Cache length: 512K.
(c) KV Cache length: 1M.
Figure 7. Computation time for single chunk prefill step on different KV Cache lengths using Qwen2-7B-Instruct. The vertical axis represents the number of attended tokens, where SDPA denotes full attention by Flashinfer (chunk size: 512).
Figure 8. End to end latency per sample with different methods on InfiniteBench using Qwen2-7B-Instruct.
6.4. Hyper-parameter Analysis
Number of selected tokens $k$.
As shown in Table 5, we fixed $n_{\text{local}}$ to a relatively small value ($512$) to compare the performance when selecting different numbers of tokens. First, we observe that even selecting a very small number of tokens (e.g., 128, 256), our TokenSelect still demonstrates very comparable performance. Then, as $k$ increases, the effectiveness of TokenSelect further improves, indicating that more moderately critical tokens also contribute to the retention of long-context information. Finally, we find that when $k$ is set to larger values (e.g., 16K), our TokenSelect shows significant improvements in most tasks, further advancing the performance landscape of long-context inference methods.
Similarity threshold of the Selection Cache $\theta$.
Fig. 6 shows that the Selection Cache hit rate increases significantly as the similarity threshold $\theta$ decreases, converging around $\theta=0.5$. This suggests potential for further acceleration of TokenSelect’s Decode Stage by reducing $\theta$. Performance sensitivity to $\theta$ varies across tasks. While most tasks exhibit slight performance degradation with decreasing $\theta$, and R.PK in InfiniteBench shows no degradation, more challenging retrieval tasks like R.KV demonstrate significant performance deterioration. This indicates higher dynamicity requirements for token selection in these tasks.
6.5. Efficiency Comparisons
Efficiency of selective sparse attention.
Fig. 7 demonstrates the significant acceleration of attention computation achieved by TokenSelect during long-context inference. With a KV Cache length of 1M, TokenSelect can provide up to 23.84×23.84$\times$23.84 × speedup compared to FlashInfer, which is the inference kernel library we based on. This substantial improvement is attributed to our efficient kernel design.
End-to-end efficiency.
Fig. 8 compares the end-to-end latency of TokenSelect, InfLLM, and standard attention across various tasks. TokenSelect significantly accelerates long-context inference in real-world scenarios, achieving a maximum speedup of 4.70×4.70$\times$4.70 × over standard attention and 2.28×2.28$\times$2.28 × over the SOTA long-context inference method. Moreover, TokenSelect demonstrates superior performance compared to both of them.
6.6. Scaling Beyond 1 Million Context Length
To further explore TokenSelect’s performance in extreme long-context scenarios, we design an extended benchmark with different text lengths following InfiniteBench. As illustrated in the Fig. 9, our TokenSelect demonstrates the ability to accurately capture critical information with a small token budget in contexts up to 2M tokens, underscoring its potential in more application scenarios.
Figure 9. Performance comparison on extended R.PK and R.KV with different text length using Qwen2-7B-Instruct.
7. Conclusion
In this paper, we introduces TokenSelect, a model-agnostic and training-free approach for efficient long-context inference and length extrapolation. TokenSelect addresses the two major challenges faced by LLMs in processing long texts: the context length limitation from pre-training and the computational complexity of attention. This is achieved through a novel token-level selective sparse attention mechanism. Experimental results demonstrate that TokenSelect can achieve up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end inference latency, while exhibiting superior performance across multiple long-context benchmarks. This approach significantly enhances LLMs’ capability to handle long contexts, paving the way for efficient long-text processing in advancing Web applications.
Appendix A Formal Statement and Proof of Lemma 1
Lemma 1 (Invariant Top-$k$ Key Selection under Cosine Similarity Threshold, Formal).
Assumptions:
  1. (1)
    Let $\mathbf{q}_{1},\mathbf{q}_{2}\in\mathbb{R}^{d}$ be two query vectors.
  2. (2)
    Let $\{\mathbf{k}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$ be a finite set of key vectors.
  3. (3)
    Let $k$ be a positive integer such that $1\leq k\leq N$.
  4. (4)
    Define the cosine similarity between vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{d}$ as:
    $\cos(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\mathbf{b}}{\|\mathbf{a}\|_{2}\|\mathbf{b}\|_{2}},$
    where $\|\cdot\|_{2}$ denotes the Euclidean norm.
  5. (5)
    Define the top-$k$ selection function based on dot product similarity as: $\mathcal{I}(\mathbf{q})=\arg\max_{S\subseteq\{1,2,\dots,N\},|S|=k}\sum_{i\in S}\mathbf{q}\cdot\mathbf{k}_{i}$. Assume that for any query vectors $\mathbf{q}$, the top-$k$ set $\mathcal{I}(\mathbf{q})$ is uniquely determined.
  6. (6)
    Let $\epsilon\in(0,1]$ be a predefined threshold.
Lemma Statement: If the cosine similarity between the two query vectors $\mathbf{q}_{1}$ and $\mathbf{q}_{2}$ satisfies
$\cos(\mathbf{q}_{1},\mathbf{q}_{2})>\epsilon,$
then the indices of the top-$k$ keys selected by $\mathbf{q}_{1}$ and $\mathbf{q}_{2}$ are identical, i.e.,
$\mathcal{I}(\mathbf{q}_{1})=\mathcal{I}(\mathbf{q}_{2}).$
Proof: We start with the given condition:
$\min_{1\leq i\leq k}{\mathbf{q}_{1}\mathbf{k}_{i}}-\max_{j>k}{\mathbf{q}_{1}\mathbf{k}_{j}}$ $>\eta,$
which we aim to use to demonstrate that:
$\min_{1\leq i\leq k}{\mathbf{q}_{2}\mathbf{k}_{i}}-\max_{j>k}{\mathbf{q}_{2}\mathbf{k}_{j}}$ $>0.$
To facilitate our analysis, we introduce the following notations:
$\hat{\eta}=\frac{\eta}{\|\mathbf{q}_{1}\|},\ \ \ \ \hat{\mathbf{q}}_{1}=\frac{\mathbf{q}_{1}}{\|\mathbf{q}_{1}\|},\ \ \ \ \hat{\mathbf{q}}_{2}=\frac{\mathbf{q}_{2}}{\|\mathbf{q}_{2}\|}.$
With these definitions, the original condition becomes:
$\min_{1\leq i\leq k}{\hat{\mathbf{q}}_{1}\mathbf{k}_{i}}-\max_{j>k}{\hat{\mathbf{q}}_{1}\mathbf{k}_{j}}$ $>\hat{\eta},$
and our goal transforms to showing:
$\min_{1\leq i\leq k}{\hat{\mathbf{q}}_{2}\mathbf{k}_{i}}-\max_{j>k}{\hat{\mathbf{q}}_{2}\mathbf{k}_{j}}$ $>0.$
Next, let $\theta$ denote the angle between $\mathbf{q}_{1}$ and $\mathbf{q}_{2}$, $\cos{\theta}=\hat{\mathbf{q}}_{1}\cdot\hat{\mathbf{q}}_{2}$. We can further define:
$\mathbf{p}_{1}=\mathbf{q}_{2}-\mathbf{q}_{1}\cos{\theta},\ \ \ \ \hat{\mathbf{p}}_{1}=\frac{\mathbf{p}_{1}}{\|\mathbf{p}_{1}\|},$
then $\sin{\theta}=\hat{\mathbf{p}}_{1}\cdot\hat{\mathbf{q}}_{2}$, and
$\hat{\mathbf{q}}_{2}=\hat{\mathbf{q}}_{1}\cos{\theta}+\hat{\mathbf{p}}_{1}\sin{\theta}.$
Then we have:
$\min_{1\leq i\leq k}{\hat{\mathbf{q}}_{2}\mathbf{k}_{i}}$ $=\min_{1\leq i\leq k}{(\hat{\mathbf{q}}_{1}\cos{\theta}+\hat{\mathbf{p}}_{1}\sin{\theta})\mathbf{k}_{i}},$
$\geq\min_{1\leq i\leq k}{\hat{\mathbf{q}}_{1}\mathbf{k}_{i}\cos{\theta}}+\min_{1\leq i\leq k}{\hat{\mathbf{p}}_{1}\mathbf{k}_{i}\sin{\theta}},$
$\geq\hat{\mathbf{q}}_{1}\mathbf{k}_{k}\cos{\theta}-\|\mathbf{k}\|_{\text{max}}\sin{\theta},$
and
$\max_{j>k}{\hat{\mathbf{q}}_{2}\mathbf{k}_{j}}$ $=\max_{j>k}{(\hat{\mathbf{q}}_{1}\cos{\theta}+\hat{\mathbf{p}}_{1}\sin{\theta})\mathbf{k}_{j}}$
$\leq\max_{j>k}{\hat{\mathbf{q}}_{1}\mathbf{k}_{i}\cos{\theta}}+\max_{j>k}{\hat{\mathbf{p}}_{1}\mathbf{k}_{i}\sin{\theta}},$
$\leq\hat{\mathbf{q}}_{1}\mathbf{k}_{p+1}\cos{\theta}+\|\mathbf{k}\|_{\text{max}}\sin{\theta}.$
Therefore,
$\min_{1\leq i\leq k}{\hat{\mathbf{q}}_{2}\mathbf{k}_{i}}$ $-\max_{j>k}{\hat{\mathbf{q}}_{2}\mathbf{k}_{j}}\geq\hat{\mathbf{q}}_{1}\mathbf{k}_{p}\cos{\theta}-\|\mathbf{k}\|_{\text{max}}\sin{\theta}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -(\hat{\mathbf{q}}_{1}\mathbf{k}_{p+1}\cos{\theta}+\|\mathbf{k}\|_{\text{max}}\sin{\theta})$
$=(\hat{\mathbf{q}}_{1}\mathbf{k}_{p}\cos{\theta}-\hat{\mathbf{q}}_{1}\mathbf{k}_{p+1}\cos{\theta})-2\|\mathbf{k}\|_{\text{max}}\sin{\theta}$
(11) $\geq\hat{\eta}\cos{\theta}-2\|\mathbf{k}\|_{\text{max}}\sin{\theta}.$
In order to have Eqn. (11) $>0$, we require
$\hat{\eta}\cos{\theta}>2\|\mathbf{k}\|_{\text{max}}\sin{\theta},$
[…]
This final inequality establishes a sufficient condition for the original statement to hold, thereby completing the proof.
Appendix B More Information on Dataset and Metrics
For InfiniteBench (Zhang et al., 2024), we use longbook_sum_eng (En.Sum), longbook_qa_eng (En.QA), longbook_choice_eng (En.MC), longdialogue_qa_eng (En.Dia), code_debug (Code.D), math_find (Math.F), passkey (R.PK), number_string (R.Num) and kv_retrieval (R.KV) as evaluation datasets. The corresponding evaluation metrics are shown in Table 6. RULER (Hsieh et al., 2024) consists of various evaluation tasks: Single NIAH (needle in a haystack), Multi-keys NIAH, Multi-values NIAH, Multi-values NIAH, Multi-queries NIAH, Variable Tracking, Common Words Extraction, Frequent Words Extraction and Question Answering. The evaluation metric is match rate. For LongBench, we use all English tasks with evaluation metrics in Table 7.
Table 6. Evaluation metrics of different datasets on InfiniteBench.
Datasets En.Sum En.QA En.MC En.Dia Code.D Math.F R.PK R.Num R.KV
Metrics Rouge-L-Sum QA F1 Score Accuracy Accuracy Accuracy Accuracy Accuracy Accuracy Accuracy
Table 7. Evaluation metrics of different datasets on LongBench.
Datasets NQA Qasper MFQA HQA 2WikiMQA Musique GovReport QMSum
Metrics QA F1 Score QA F1 Score QA F1 Score QA F1 Score QA F1 Score QA F1 Score Rouge-L Rouge-L
Datasets MultiNews TREC TQA SAMSum PsgCount PsgRetrieval LCC RepoBench-P
Metrics Rouge-L Accuracy QA F1 Score Rouge-L Accuracy Accuracy Code Sim Score Code Sim Score
Table 8. Comparison of different methods with different origin models on LongBench.
Methods NQA Qasper MFQA HQA 2WikiMQA Musique GovReport QMSum MultiNews
Qwen2-7B 24.24 45.42 47.79 42.76 44.38 24.16 33.80 23.78 26.17
NTK 26.25 45.94 50.76 53.20 50.31 30.83 32.75 23.21 25.94
SelfExtend 7.15 20.37 24.06 14.91 13.73 4.75 16.92 16.53 18.74
StreamLLM 19.49 42.56 39.63 42.43 44.67 15.22 31.51 20.57 26.00
InfLLM 27.47 41.44 46.99 47.47 49.29 25.62 32.68 23.10 26.77
TokenSelect 24.18 42.29 45.77 48.62 49.08 27.85 33.69 23.03 26.35
Llama-3-8B 19.85 42.36 41.03 47.38 39.20 22.96 29.94 21.45 27.51
NTK 9.90 45.35 49.41 48.86 29.22 24.56 34.31 23.82 27.27
SelfExtend 1.72 8.90 20.80 8.65 6.97 3.27 13.99 15.36 17.66
StreamLLM 20.05 42.46 39.54 43.69 37.89 19.68 29.17 21.33 27.56
InfLLM 22.64 43.70 49.03 49.04 35.61 26.06 30.76 22.70 27.57
TokenSelect 22.44 40.74 47.73 50.33 31.38 24.53 32.56 23.50 27.92
Yi-1.5-6B 17.18 32.56 39.06 36.26 39.25 16.32 30.53 20.21 26.20
NTK 0.80 35.06 29.05 7.47 24.38 0.73 13.66 6.25 25.43
SelfExtend 3.29 19.03 26.00 17.11 11.88 7.73 20.38 17.46 21.79
StreamLLM 15.05 33.27 38.31 34.91 36.92 16.33 29.38 20.02 26.14
InfLLM 17.65 36.25 45.40 41.25 35.89 16.94 30.22 20.85 26.04
TokenSelect 19.36 33.98 48.14 45.05 40.13 22.98 31.59 21.51 26.48
Methods TREC TQA SAMSum PsgCount PsgRetrieval LCC RepoBench-P Average
Qwen2-7B 78.50 88.77 46.33 5.50 70.00 62.40 61.95 45.37
NTK 79.50 89.51 46.03 5.50 60.00 59.36 59.69 46.17
SelfExtend 16.50 27.54 29.42 4.50 0.00 41.42 41.89 18.65
StreamLLM 75.50 87.19 46.27 3.50 27.50 61.18 61.12 40.27
InfLLM 70.50 87.51 44.53 4.00 46.50 55.08 57.53 42.90
TokenSelect 74.00 89.26 45.94 5.00 42.50 61.48 59.33 43.64
Llama-3-8B 74.00 90.50 42.30 8.50 62.50 60.83 49.14 42.46
NTK 73.00 88.74 42.51 8.87 99.50 33.62 35.04 42.12
SelfExtend 20.50 16.82 25.39 5.75 7.50 26.24 31.22 14.42
StreamLLM 73.50 90.08 41.55 5.00 49.00 60.35 48.95 40.61
InfLLM 73.50 90.91 42.43 7.17 84.00 59.88 46.48 44.46
TokenSelect 67.50 92.22 42.16 4.54 87.00 58.86 51.24 44.04
Yi-1.5-6B 71.50 48.79 0.79 3.00 28.50 57.10 52.53 32.48
NTK 40.00 12.71 1.34 0.50 3.35 54.55 37.24 18.28
SelfExtend 23.75 30.61 2.58 2.75 13.50 43.17 35.45 18.53
StreamLLM 69.00 73.36 0.82 2.50 18.50 56.37 49.05 32.49
InfLLM 71.50 71.49 1.01 4.00 10.50 56.88 46.28 33.25
TokenSelect 62.50 69.70 0.62 3.50 41.50 54.32 54.99 36.02
Appendix C Experimental Results on LongBench
Compared to InfiniteBench and RULER, LongBench has much shorter text lengths. The 95% percentile for its lengths is 31K tokens. Considering that recent LLMs after SFT generally have context lengths of up to 32K tokens (Yang et al., 2024), LongBench is less suitable for evaluating state-of-the-art long-context inference methods. Nevertheless, as shown in Table 8, our TokenSelect still demonstrates superior overall performance compared to most baseline methods. It’s worth noting that Yi-1.5-6B did not yield effective results on the SAMSum task because it failed to correctly follow instructions.